Grade 12Magical Mathematics[Interesting Approach]

In throwing of a pair of dice find the probability of the event total is not eat and not 11

Manish patel
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear Manish
 
sum is 8 = (6,2)(2,6)(5,3)(3,5)(4,4)
hence probability of getting 8 = 5 / 36
sum is 11 = (6,5)(5,6)
hence probability of getting 11 = 2 / 36
now 
P (A U B) = 5 + 2 / 36 = 7 / 36
Hence nether 8 come nor 11, probability = P (A U B) ‘ =  1 –  7 / 36 = 29 / 36
 
Regards
Arun (askIITians forum expert)